OpenStudy (anonymous):

Find the result of : 1^3 - 2^3 + 3^3 - 4^3 + ... + 49^3 - 50^3

OpenStudy (anonymous):

do u have formula or property or something to simplify it

OpenStudy (anonymous):

1^3 - 2^3 + 3^3 - 4^3 + ... + 49^3 - 50^3 =(1-2)(1^2 + 1*2+2^2) + (3-4) (3^2 + 3*4 + 4^2) + ... + (49-50)(49^2 + 49*50+50^2) = - (1^2 + 2^2 + 3^2 + 4^2 +... + 50^2) - ( 1*2 + 3*4 + 5*6 + ... + 49*50)

OpenStudy (anonymous):

how do u get that, what formula have u used ?

OpenStudy (anonymous):

oh, i see. from. (a-b)^3 = a^3 - 3a^2b + 3ab^2 - b^3 so, a^3 - b^3 = (a-b)^3 + 3ab(a-b) right ??

OpenStudy (anonymous):

but i dont see like u done it, what formula is it ?

OpenStudy (anonymous):

a^3 - b^3 = (a-b) (a^2 + ab + b^2)

OpenStudy (anonymous):

oh yea, you are correct. next, how do u calculate of 1^2 + 2^2 + 3^2 + 4^2 +... + 50^2 ? similar question for 1*2 + 3*4 + 5*6 + ... + 49*50) ?

OpenStudy (anonymous):

1^2 + 2^2 + 3^2 + 4^2 +... + 50^2 is calculated by sum of n squares .. there is a formula : n(n+1)(2n+1) / 6 plug in it 50.

OpenStudy (anonymous):

ok, thanks.. so, 1^2 + 2^2 + 3^2 + 4^2 +... + 50^2 = 50(51)(101)/6 = 42,925

OpenStudy (anonymous):

how about u with 1*2 + 3*4 + 5*6 + ... + 49*50) ?

OpenStudy (anonymous):

now for 1*2 + 3*4 + 5*6 + ... + 49*50 we have for the nth term n(n+1) = n^2 + n so we have again something like 1^2 + 1 + 2^2 + 2 + 3^2 + 3 + ... so we have again sum of n squares but this time the last is 49 and we have sum of n (not squares ) which is given by n(n+1)/2 again last term is 49

OpenStudy (anonymous):

so, i see that 1^2 + 1 + 2^2 + 2 + 3^2 + 3 + ... = (1^2 + 2^2 + 3^2 + ...+ 49^2) + (1+2+3+ ... + 49) right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, now make sense for me... thank u very much @sauravshakya and @Coolsector

OpenStudy (anonymous):

:)